OpenStudy (anonymous):

A ray of light is traveling from water into a diamond. Upon hitting the diamond, a light ray bends toward the normal. With this information, what can you say about the speed of light as it passes from water into a diamond?

OpenStudy (anonymous):

a. speed increases b. speed decreases c.speed drops to 0 d.speed doesnt change

OpenStudy (theeric):

I have to go, but really quick!That might help.. Think about speed.... Is light faster in more or less dense materials? The fact that it bent like that, does that mean diamond is more or less dense? You can also visualize it, but I don't have time to explain that right now. Good luck!

OpenStudy (anonymous):

n = Index of refraction = speed of light in vacuum / speed of light in medium

OpenStudy (theeric):

You can be sure that it's not C or D. Light doesn't stop, and it's change in speed is the entire reason for refraction! Down to two. in addition, there's Snell's Law. http://en.wikipedia.org/wiki/Snell%27s_law \(\dfrac{\sin\theta_1}{\sin\theta_2} = \dfrac{v_1}{v_2} = \dfrac{n_2}{n_1}\) The angles, \(\theta_1\) and \(\theta_2\) are measured from the surface normal on their side of the medium. This picture shows what I mean: http://upload.wikimedia.org/wikipedia/commons/thumb/3/3f/Snells_law2.svg/220px-Snells_law2.svg.png That's a lot right there. Let's look at the first two of the equivalent expressions. \(\dfrac{\sin\theta_1}{\sin\theta_2} = \dfrac{v_1}{v_2} \) Because we see which angle is bigger, we can know where the velocity is bigger. If you think of the sine function, an angle of zero gives you a ____... And as it get's larger, the sine of that angle gets _______.... Combine that with the fact that if the sine on top is larger than the bottom, then the velocity on top is bigger than the bottom - and vise versa.

OpenStudy (theeric):

Feel free to ask if you have further questions!